\begin{thebibliography}{99} % Updated count to N unique references

    \bibitem{NayakRMP} Nayak, C., Simon, S. H., Stern, A., Freedman, M., \& Das Sarma, S. (2008). Non-Abelian anyons and topological quantum computation. \textit{Reviews of Modern Physics}, 80(3), 1083.

    \bibitem{FreedmanLarsenWang} Freedman, M., Larsen, M., \& Wang, Z. (2002). A modular functor which is universal for quantum computation. \textit{Communications in Mathematical Physics}, 227, 605--622.

    \bibitem{Feiguin2007} Feiguin, A.; Trebst, S.; Ludwig, A. W. W.; Troyer, M.; Kitaev, A.; Wang, Z.; Freedman, M. H. Interacting Anyons in Topological Quantum Liquids: The Golden Chain. \textit{Physical Review Letters} \textbf{2007}, 98, 160409.

    \bibitem{IqbalNature2024} Iqbal, M., Tantivasadakarn, N., Verresen, R. et al. Non-Abelian topological order and anyons on a trapped-ion processor. \textit{Nature} \textbf{626}, 505–511 (2024). 

    \bibitem{Amaral2022} Amaral, M.; Chester, D.; Fang, F.; Irwin, K. Exploiting Anyonic Behavior of Quasicrystals for Topological Quantum Computing. \textit{Symmetry} \textbf{2022}, 14, 1780. 

    \bibitem{LS86} Levine, D.; Steinhardt, P. J. Quasicrystals. I. Definition and structure. \textit{Phys. Rev. B} \textbf{1986}, 34, 596-616. 

    \bibitem{BaakeGrimmBook} Baake, M.; Grimm, U. \textit{Aperiodic Order: Volume 1, A Mathematical Invitation}; Cambridge University Press, 2013. 

    \bibitem{TemperleyLieb} Temperley, H. N. V.; Lieb, E. H. Relations between the 'percolation' and 'colouring' problem and other graph-theoretical problems associated with regular planar lattices: some exact results for the 'percolation' problem. \textit{Proc. R. Soc. Lond. A} \textbf{1971}, 322(1549), 251-280.

    \bibitem{KauffmanLomonaco} Kauffman, L. H.; Lomonaco Jr., S. J. The Fibonacci Model and the Temperley-Lieb Algebra. \textit{Int. J. Mod. Phys. B} \textbf{2008}, 22(29), 5065-5087. 

    \bibitem{Qiskit}
    A.\ Javadi-Abhari, M.\ Treinish, K.\ Krsulich, C.\ J.\ Wood, J.\ Lishman, J.\ Gacon, S.\ Martiel, P.\ D.\ Nation, L.\ S.\ Bishop, A.\ W.\ Cross, B.\ R.\ Johnson, and J.\ M.\ Gambetta, ``Quantum computing with \textsc{Qiskit},''
    \emph{arXiv}:2405.08810 [quant-ph] (2024).
    
    \bibitem{QIC_core_GitHub}
    M.~M.~Amaral, ``quasicrystal-inflation-code,'' GitHub repository,\\
    \url{https://github.com/macielamaral/quasicrystal-inflation-code}, May 2025.

    \bibitem{Laakkonen2025}
    T.~Laakkonen, E.~Rinaldi, C.~N.~Self, E.~Chertkov, M.~DeCross, D.~Hayes, B.~Neyenhuis, M.~Benedetti, K.~Meichanetzidis,
    ``Less Quantum, More Advantage: An End-to-End Quantum Algorithm for the Jones Polynomial,''  
    \emph{arXiv:2503.05625} [quant-ph], 2025.
 
    \bibitem{Jones1990}
    V.~F.~R.~Jones, 
    ``A polynomial invariant for knots via von Neumann algebras,'' 
    in \emph{New Developments in the Theory of Knots}, 
    Advanced Series in Mathematical Physics, vol.~11, pp.~3--11, 1990.

    \bibitem{KnotInfo-12a122}
    C. Livingston and A. H. Moore, KnotInfo: Table of Knot Invariants, knotinfo.math.indiana.edu, May 01, 2025. 
    
    \bibitem{AmaralHierarchy2025}
    Amaral, M., Aschheim, R. (2025). 
    From Forbidden Words to Anyon Models: A General Hamiltonian from a Constraint Hierarchy. (Work in preparation).

\end{thebibliography}

\documentclass[11pt, a4paper]{article} 
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{geometry}
\usepackage[utf8]{inputenc}
\usepackage{txfonts}
\usepackage[english]{babel}
\usepackage{cite} 
\usepackage{hyperref}

\newcommand{\phiGR}{\phi}
\newcommand{\ket}[1]{\left| #1 \right\rangle}
\newcommand{\bra}[1]{\left\langle #1 \right|}
\newcommand{\Hamiltonian}{H_{QIC}} % Symbol for Hamiltonian operator/matrix
\newcommand{\HilbertQIC}{\mathcal{H}_{QIC}} % Symbol for QIC subspace
 
\newcommand{\Identity}{I} % Identity operator

\title{\textbf{Consistent Simulation of Fibonacci Anyon Braiding\\ within a Qubit Quasicrystal Inflation Code}}
\author{Marcelo M. Amaral\\
\textit{\small{Quantum Gravity Research – Los Angeles}}\\
\textit{\small{AperiodiQ – Los Angeles}}
}
\date{\today}

\begin{document}
\maketitle

\begin{abstract}
The simulation of non-Abelian anyon braiding is a critical step towards fault-tolerant quantum computation. We introduce a framework for this task based on a one-dimensional \textit{Quasicrystal Inflation Code} (QIC). The code is defined by a local Hamiltonian whose ground state manifold enforces Fibonacci tiling constraints and possesses the correct Fibonacci degeneracy. We derive the corresponding braid operators and demonstrate that, while they are formally non-local, they possess an exact, local 3-qubit structure. This allows us to distill their action into a single, physically realizable $8 \times 8$ gate, which we term the $B_{gate}$.
We prove through comparative compilation for systems with different numbers of qubits that constructing circuits by composing these local gates is dramatically more scalable than compiling the equivalent global unitary, showing a greater than tenfold reduction in circuit depth. To validate the framework, we successfully executed a braiding algorithm for the Jones polynomial, a topological invariant of knots, on an IBM Quantum processor, quantifying the signal degradation due to hardware noise. 
Finally, we numerically confirm for systems up to 17 qubits that our construction rigorously satisfies the required Temperley–Lieb and braid group relations while preserving the code subspace. This work establishes a validated and physically grounded pathway for the scalable simulation of anyonic braiding on programmable quantum systems.
\end{abstract}

\section{Introduction} 
\label{sec:intro}

Topological quantum computation (TQC) offers a compelling route towards fault-tolerant quantum information processing by encoding information non-locally within the states of topological systems \cite{NayakRMP}. Systems supporting non-Abelian anyons, particularly Fibonacci anyons, are of great interest as their braiding operations are sufficient for universal quantum computation \cite{FreedmanLarsenWang,Feiguin2007}. However, the direct physical realization and manipulation of anyons remain significant experimental hurdles across various proposed platforms.

An alternative strategy involves simulating the essential physics of anyonic systems, including their characteristic fusion rules and braiding statistics, on more readily available quantum computing architectures, such as superconducting qubits or trapped ions. Recent experiments have successfully demonstrated the creation and control of non-Abelian topological order and anyons on such platforms, showcasing the potential and challenges of this approach \cite{IqbalNature2024}. This requires robust encoding schemes that faithfully map the abstract algebraic structure of anyon models onto the qubit Hilbert space while ideally providing some level of inherent protection or structure.

Quasicrystal Inflation Codes (QIC) represent a promising framework for this purpose, drawing inspiration from the structural properties of quasicrystals and aperiodic tilings \cite{Amaral2022}. This work investigates the Fibonacci QIC realized on a one-dimensional (1D) chain of qubits. We leverage the structure of 1D Fibonacci quasicrystals, whose physical generation via tiling growth or substitution rules (e.g., mapping Long tiles L to $\ket{1}$ and Short tiles S to $\ket{0}$) naturally forbids adjacent identical short tiles ('SS') \cite{LS86, BaakeGrimmBook}. Indeed, the iterative application of these substitution rules, effectively mapping the quasicrystal structure from one length scale to the next, ensures that S tiles (representing $\ket{0}$) either transform into L tiles ($\ket{1}$) or are newly generated only as part of an LS sequence derived from a parent L tile. This process inherently prevents the formation of 'SS' (or $\ket{00}$) configurations. This fundamental physical constraint directly motivates the definition of the QIC via the exclusion of adjacent $\ket{00}$ states in the qubit chain.

We define the protected logical subspace of this standard Fibonacci QIC, denoted $\HilbertQIC$, as the subspace of the $N$-qubit Hilbert space $\mathcal{H}$ spanned by all computational basis states satisfying the 'no 00' constraint. To make this concrete, we introduce a local Hamiltonian $\Hamiltonian$ (detailed in Sec.~\ref{sec:hamiltonian}, Eq.~\eqref{eq:H_QIC_rep}) which energetically penalizes $\ket{00}$ configurations, ensuring $\HilbertQIC$ is its zero-energy ground state manifold. Crucially, the dimension of this physically motivated subspace is $\dim(\HilbertQIC) = F_{N+2}$ (where $F_n$ is the $n$-th Fibonacci number). This dimension is directly comparable to that of the fusion space for $N$ abstract Fibonacci anyons; indeed, Hilbert space dimensions for such anyonic systems, such as $F_{N+1}$ or $F_{N+2}$, are known to depend on the specific model conventions and boundary conditions employed \cite{Feiguin2007, NayakRMP}. This establishes $\HilbertQIC$ as a valid encoding space for a standard Fibonacci anyon model requiring $F_{N+2}$ dimensions.

Simulating TQC within $\HilbertQIC$, however, requires implementing operators that correctly represent anyon braiding. These operations are mathematically described by the Temperley–Lieb (TL) algebra $TL_N(\delta)$ with parameter $\delta = \phiGR$ (where $\phiGR$ is the golden ratio)~\cite{TemperleyLieb,KauffmanLomonaco}. A challenge arises when attempting to realize this algebra on qubits within constrained subspaces. While models like the Golden Chain~\cite{Feiguin2007} define local projectors satisfying the TL algebra in the anyon fusion basis, directly adapting such local constructions to act on qubits within $\HilbertQIC$ does not automatically ensure the crucial multi-site TL relations (Eqs.~\eqref{eq:Pn_TL_rep2}--\eqref{eq:Pn_commute_rep2}) are preserved. 

This work addresses the critical implementation gap by demonstrating that a complete set of braid operations can be defined and validated on the full qubit Hilbert space. We present and rigorously verify a method to consistently represent and simulate the $TL_N(\phiGR)$ algebra and its associated braid group generators within the Hamiltonian-defined QIC subspace $\HilbertQIC$. Our approach proceeds in two steps: (i) we construct the abstract $F_{N+2}$-dimensional Temperley–Lieb projector matrices $P_n$, which act correctly on the QIC basis states in accordance with established diagrammatic rules~\cite{KauffmanLomonaco}; and (ii) we embed these operators into the full $2^N$-dimensional qubit Hilbert space via a numerically derived isometry $V$, yielding $P'_n = V P_n V^\dagger$ (see Sec.~\ref{sec:embedding}). This embedding ensures that the algebraic structure of the anyonic model is preserved at the physical qubit level.

Through extensive numerical simulations performed for systems up to $N=17$ qubits (Sec.~\ref{sec:nonlocal_impl_verif}), we demonstrate that these embedded operators $P'_n$ rigorously satisfy the defining relations of the $TL_N(\phiGR)$ algebra (Eqs.~\eqref{eq:Pn_projector_rep2}--\eqref{eq:Pn_commute_rep2}). Consequently, the braid operators $B'_n$ constructed from them (Eq.~\eqref{eq:Bn_def_sim_rep2}) are unitary, satisfy the Yang–Baxter equation and required commutation relations, and demonstrably preserve the protected QIC subspace $\HilbertQIC$ during operation. This provides a concrete and numerically validated framework for simulating Fibonacci anyon TQC on standard qubit platforms, leveraging physically motivated quasicrystal constraints.

While the non-local isometric embedding faithfully encodes the full braid structure within the QIC subspace, its global nature renders it impractical for near-term quantum hardware implementations. To overcome this limitation, we extract a local, physically realizable form of the braid operations (Sec.~\ref{sec:local_bgate}), $B_{gate}$. Notably, each operator $B_n'$ can be represented exactly by a unitary acting on just three qubits. This inherent locality enables direct quantum simulation within the code space and, as shown in our transpilation study (Sec.~\ref{sec:scalable_compilation}), leads to a substantial reduction in circuit depth---achieving over an order of magnitude improvement compared to brute-force compilation of the global unitary. We demonstrate the effectiveness of this framework through the computation of Jones polynomials (Sec.~\ref{sec:jones_application}), a Qiskit-based \cite{Qiskit} simulation for the $N=3$ case (Sec.~\ref{sec:qiskit_sim}), hardware performance analysis (Sec.~\ref{sec:hardware_execution}), and a comparison with alternative QIC models (Sec.~\ref{sec:relation_to_hierarchy}).

This work addresses a critical gap between abstract anyonic models and physical quantum hardware by presenting a complete framework from physical encoding to scalable implementation. Our approach involves: (i) defining the protected code space as the ground state of a physically motivated local Hamiltonian, $\Hamiltonian$; (ii) constructing abstract Temperley–Lieb operators and embedding them via an isometry to obtain theoretically correct braid generators, $B'_n$; and (iii) discovering that these operators can be realized as local, efficient 3-qubit gates acting directly within the code space. Through extensive simulations and a proof-of-principle hardware demonstration, we validate that the proposed approach offers a scalable and physically grounded route to simulate anyonic braiding on standard quantum computing platforms.

\section{The QIC Hamiltonian and Protected Subspace} 
\label{sec:hamiltonian}

We consider a system of $N$ qubits with the total Hilbert space $\mathcal{H} = (\mathbb{C}^2)^{\otimes N}$. 
The QIC Hamiltonian we introduce is derived from the fundamental structural properties of one-dimensional (1D) Fibonacci quasicrystalline tilings or sequences (we note that also apply to 2D quasicrystals made of two tiles). The generation of such physical systems is governed by well-established substitution rules, a canonical example being $L \to LS$ and $S \to L$ (where L denotes a Long tile and S a Short tile). A direct consequence of these rules is the inherent prohibition of configurations containing adjacent identical "short" tiles (denoted 'SS') \cite{LS86, BaakeGrimmBook, Amaral2022}. By mapping the short tile S to the qubit state $\ket{0}$ and the long tile L to $\ket{1}$, this physical constraint translates directly into the defining rule for our QIC: the exclusion of adjacent $\ket{00}$ states in the $N$-qubit chain.

Our Hamiltonian, $\Hamiltonian$, is designed to energetically enforce exactly this physically motivated constraint. Let $Z_i$ be the Pauli-Z operator on qubit $i$. The local projector onto the forbidden $\ket{00}$ state at adjacent sites $i, i+1$ is given by $\Pi^{(00)}_{i+1, i} = \frac{1}{4}(\Identity + Z_{i+1})(\Identity + Z_i)$. The QIC Hamiltonian is then defined as a sum of these local penalty terms:
\begin{equation}
    \Hamiltonian = \lambda \sum_{i=0}^{N-2} \Pi^{(00)}_{i+1, i} \quad (\lambda > 0)
    \label{eq:H_QIC_rep}
\end{equation}
The ground state manifold of $\Hamiltonian$ (with energy 0) defines the QIC subspace, denoted $\HilbertQIC \subset \mathcal{H}$. By construction, $\HilbertQIC$ is spanned by the set of all computational basis states $\ket{s_{N-1} \dots s_0}$ of length $N$ that contain no adjacent '00' substring.

The dimension of this subspace follows the Fibonacci recurrence and is given by $\dim(\HilbertQIC) = F_{N+2}$, where $F_n$ are the standard Fibonacci numbers with $F_0=0, F_1=1$ (sequence 0, 1, 1, 2, 3, 5, 8, ...). This dimensionality is the well-known result for binary sequences avoiding '00' and, critically, precisely matches the dimension of the fusion space of $N$ Fibonacci anyons \cite{Feiguin2007}. An orthonormal basis $\{ \ket{q_a} \}_{a=1}^{F_{N+2}}$ for $\HilbertQIC$, corresponding for instance to the lexicographically sorted valid 'no 00' binary strings, can be algorithmically constructed.

Thus, this explicit Hamiltonian $\Hamiltonian$ provides a concrete qubit realization whose protected code space $\HilbertQIC$ is directly motivated by physical Fibonacci quasicrystal constraints and possesses the correct dimensionality required for simulating the standard $N$-anyon Fibonacci model.

\section{Required Anyon Operators and Algebra}
\label{sec:operators}

With the protected subspace $\mathcal{H}_{QIC}$ established, performing quantum computation requires a set of logical operators that simulate anyon braiding. This physical process is governed by the Temperley-Lieb (TL) algebra, $TL_N(\delta)$ \cite{TemperleyLieb, KauffmanLomonaco}, with the parameter for the Fibonacci model set to the golden ratio, $\delta = \phiGR = (1+\sqrt{5})/2$. The generators of this algebra are a set of Temperley-Lieb projectors, $P_n$. For a system of size $N$, these projectors generate a specific instance of the algebra, $\mathcal{A}_N = \mathrm{Alg}(P_0, \dots, P_{N-2})$. These algebras form a nested structure, $\mathcal{A}_N \subset \mathcal{A}_{N+1}$, which is analogous to the Jones tower construction in operator algebras.

\subsection{Temperley-Lieb Projectors ($P_n$)}
The generators of the algebra $\mathcal{A}_N$ are the TL projectors $P_n$ ($n=0, \dots, N-2$). Abstractly, these operators act on the $F_{N+2}$-dimensional QIC basis $\HilbertQIC$ (equivalent to the $N$-anyon fusion basis or the $N$-level inflation basis in a growing quasicrystal \cite{Amaral2022}) and satisfy the following defining relations:
\begin{align}
    P_n^2 &= P_n && \text{(Idempotent)} \label{eq:Pn_projector_rep2}\\
    P_n^\dagger &= P_n && \text{(Hermitian)} \label{eq:Pn_hermitian_rep2}\\
    P_n P_{n\pm 1} P_n &= \phiGR^{-2} P_n && \text{(TL Relation)} \label{eq:Pn_TL_rep2}\\
    P_n P_m &= P_m P_n && \text{for } |n-m| \ge 2 \label{eq:Pn_commute_rep2}
\end{align}
%Physically, within the anyon model, the projector $P_n$ acts locally on anyons at positions $n$ and $n+1$, projecting onto the identity (vacuum) fusion channel $\Identity$, while $(\Identity-P_n)$ projects onto the Fibonacci anyon ($\tau$) channel. 
Note that Kauffman-Lomonaco in \cite{KauffmanLomonaco} uses the operator $U_n$, which is related to $P_n$ by $U_n=\phiGR P_n$. 

\subsection{Braid Operators ($B_n$)}
Unitary operators $B_n$ representing the physical process of braiding adjacent anyons at positions $n$ and $n+1$ are constructed from the TL projectors $P_n$ (in (Sec.~\ref{sec:relation_to_hierarchy}) we will comment on the physical scenario within quasicrystals). Here, we adapt the notation from \cite{KauffmanLomonaco, Amaral2022}. This yields the braid operator:
\begin{equation}
    B_n = R_I P_n + R_{\tau} (\Identity - P_n) = e^{-4\pi i/5} P_n + e^{3\pi i/5} (\Identity - P_n)
    \label{eq:Bn_def_sim_rep2}
\end{equation}
where $\Identity$ is the identity operator on the $F_{N+2}$-dimensional space, and $R_I = e^{4\pi i/5}$ and $R_{\tau} = e^{-3\pi i/5}$ are the topological phases (R-matrix eigenvalues) associated with the identity and $\tau$ fusion channels, respectively. These operators $B_n$ satisfy the Braid Group relations required for valid TQC operations (including the Yang-Baxter equation $B_n B_{n+1} B_n = B_{n+1} B_n B_{n+1}$ and the commutation relation $B_n B_m = B_m B_n$ for $|n-m| \ge 2$) if and only if the underlying projectors $P_n$ satisfy the TL algebra relations Eqs.~\eqref{eq:Pn_projector_rep2}-\eqref{eq:Pn_commute_rep2}.

From the definition Eq.~\eqref{eq:Bn_def_sim_rep2}, one can directly verify the consistency relation:
\begin{equation}
    P_n B_n = B_n P_n = e^{4\pi i/5} P_n. \label{eq:PB_consistency}
\end{equation}
This confirms that $P_n$ projects onto the subspace (the identity fusion channel) where the braid operator $B_n$ acts simply as multiplication by the phase $R_I = e^{4\pi i/5}$. 

Note that Eq.~\eqref{eq:Bn_def_sim_rep2} is equivalent to the braid representation definition from \cite{Amaral2022}, which can be seeing by writing $\phiGR A = R_I - R_{\tau}$, where $A=e^{3\pi i/5}$. And our braid and its inverse definition is switched from \cite{KauffmanLomonaco}.

\section{Embedding into Qubit Space}
\label{sec:embedding}

To simulate these operations on qubits, we embed the abstract operators into the full $2^N$-dimensional Hilbert space $\mathcal{H}$ while preserving their action within the QIC subspace $\mathcal{H}_{QIC}$.
\begin{enumerate}
    \item \textbf{Isometry ($V$):} We construct the isometry $V: \mathbb{C}^{F_{N+2}} \to \HilbertQIC \subset \mathcal{H}$ mapping the abstract $F_{N+2}$-dimensional basis (ordered by QIC strings) to the corresponding orthonormal state vectors $\{\ket{q_a}\}$ within $\mathcal{H}$. $V$ satisfies $V^\dagger V = I_{F_{N+2}}$.
    \item \textbf{Embedded Operators ($P'_n, B'_n$):} The abstract operators $P_n$ and $B_n$ (acting on $\mathbb{C}^{F_{N+2}}$) are embedded into $\mathcal{H}$ as $2^N \times 2^N$ matrices:
        \begin{align}
            P'_n &= V P_n V^\dagger \\
            B'_n &= V B_n V^\dagger
        \end{align}
\end{enumerate}
These operators $P'_n$ and $B'_n$ act non-trivially only within the QIC subspace $\HilbertQIC$. This is because they are constructed such that $P'_n = P'_n \Pi_{QIC} = \Pi_{QIC} P'_n$ (and similarly for $B'_n$), where $\Pi_{QIC} = VV^\dagger$ is the projector onto $\HilbertQIC$. Consequently, if the operators $P_n$ satisfy the Temperley-Lieb algebra, the embedded $2^N \times 2^N$ matrices $P'_n$ satisfy the same algebraic relations on the full Hilbert space $\mathcal{H}$. This algebraic fidelity ensures that their action on states within $\HilbertQIC$ correctly simulates the anyonic system. Similarly, the derived $B'_n$ operators satisfy the required braid group relations (such as the Yang-Baxter equation) and, by construction, preserve the $\HilbertQIC$ subspace. Note that $B'_n$ as defined by $V B_n V^\dagger$ is equivalent to $B'_n = R_I P'_n + R_{\tau} (\Pi_{QIC} - P'_n)$.

%--------------------------------------------------
\subsection{From a Global Operator to a Local 3-Qubit Gate}
\label{sec:local_bgate}

While the embedded braid operators $B'_n = V B_n V^\dagger$ are formally $2^N \times 2^N$ matrices acting on the full qubit Hilbert space, a critical analysis reveals a simplification. The action of a given generator $B'_n$, which represents a braid between strands $n$ and $n+1$, is non-trivial only on the computational basis states involving the corresponding physical qubits $n{-}1$, $n$, and $n{+}1$. The operator acts as the identity on all other ``spectator'' qubits.

This inherent locality allows us to distill the operator's action down to a single, concrete $8 \times 8$ unitary matrix, which we term the local $B_{gate}$. This is not an approximation, but an exact, physically realizable building block for braiding, which can be applied to any triplet of adjacent qubits $(n{-}1,n,n{+}1)$ within a larger quantum circuit. The $B_{gate}$ is constructed from a corresponding local projector $P_{\text{local}}$ as follows:
\begin{equation}
B_{gate} = R_I P_{\text{local}} + R_{\tau} (\mathbb{I} - P_{\text{local}})
\end{equation}
where $R_I$ and $R_{\tau}$ are the eigenvalues associated with the braid generator in the Fibonacci model. The explicit form of $B_{gate}$, as generated by our simulation code, is provided in Appendix~\ref{sec:canonicalBgateExp}.

The construction of $P_{\text{local}}$ proceeds by analyzing its action on the 8-dimensional Hilbert space of three qubits. The QIC constraint---which forbids adjacent $\ket{00}$ configurations---partitions this space into a five-dimensional valid subspace,
\[
\mathcal{H}_{\text{valid}} = \mathrm{span}\{\ket{010}, \ket{011}, \ket{101}, \ket{110}, \ket{111}\},
\]
and a three-dimensional forbidden subspace. We first use the Kauffman–Lomonaco diagrammatic rules to construct the $5 \times 5$ matrix describing the projector on $\mathcal{H}_{\text{valid}}$. This is then embedded into an $8 \times 8$ matrix $P_{\text{local}}$ by defining it to act as the identity on the forbidden subspace. 

This construction guarantees two essential properties. First, since $P_{\text{local}}$ is Hermitian and idempotent by construction, the resulting $B_{gate}$ is manifestly unitary. Second—and most critically—the $B_{gate}$ preserves the QIC subspace: for any state $\ket{\psi}$ in $\mathcal{H}_{\text{valid}}$, the transformed state $B_{gate} \ket{\psi}$ is also confined to $\mathcal{H}_{\text{valid}}$. States in the forbidden subspace acquire only a global phase, leaving the code unaffected. This confirms that $B_{gate}$ is a legal operation within our QIC framework, fully compatible with its constraints.

The existence of this local gate provides a major scalability advantage. A braid word of arbitrary length can be implemented as a sequence of these efficient 3-qubit gates, yielding a circuit depth that grows linearly with the number of braids. This stands in stark contrast to the ``global unitary'' strategy, where compiling the full $2^N \times 2^N$ operator becomes infeasible due to exponential scaling. As demonstrated in our transpilation study (Sec.~\ref{sec:local_impl_verif}), the local approach yields over an order-of-magnitude reduction in circuit depth for $N=6$. While the existence of such local gates is known in the TQC literature, our contribution lies in rigorously deriving it from first principles within the QIC formalism, proving its exact equivalence to $B'_n$, and quantifying its practical advantage for compilation on real quantum hardware.

It is critical to distinguish this local, translationally-invariant $B_{gate}$ from the operators generated by the non-local embedding pipeline for a finite chain. For a small system such as $N=3$, the non-local construction correctly yields operators (e.g., $B'_0, B'_1$) that are specific to that finite size and include inherent boundary effects, making them non-identical. To obtain the canonical, reusable bulk gate, we developed a set of direct construction functions \cite{QIC_core_GitHub} that build the ideal $B_{gate}$ independent of system size. This gate is the fundamental building block for all scalable braiding circuits, representing the true bulk physics of the anyon interaction.

During the completion of this work, we noted the concurrent release of a similar hardware implementation for Jones polynomial evaluation by Laakkonen et al. \cite{Laakkonen2025}. Their work also arrives at a local 3-qubit gate for the Fibonacci model, which is fundamentally equivalent to our \texttt{B\_gate} as both are ultimately derived from the Kauffman-Lomonaco formalism \cite{KauffmanLomonaco}. The specific convention chosen in our framework (Sec.~\ref{sec:operators}) defines our braid operator as the inverse of that used in \cite{KauffmanLomonaco}, which consequently makes our \texttt{B\_gate} equivalent to the inverse of the gate presented in \cite{Laakkonen2025}. This physical equivalence is benchmarked in Sec.~\ref{sec:jones_application}, where the results of our framework are compared against theirs for the calculation of the Jones polynomial.

\section{QIC Non-Local Implementation and Verification}
\label{sec:nonlocal_impl_verif}

To correctly simulate anyon braiding on a qubit substrate, the abstract TL algebra and Braid Group operators must be faithfully represented within the full qubit Hilbert space, $\mathcal{H}$. This section details the successful construction and verification of these operators by embedding them into the qubit space via a computationally constructed isometry. This "non-local" approach, while not scalable for direct circuit implementation, is a crucial step to prove the theoretical validity of the QIC framework and serves as the foundation from which practical, local gates are derived.

The core of our method is a pipeline that translates the abstract $F_{N+2}$-dimensional anyonic operators into concrete $2^N \times 2^N$ matrices acting on the qubits. 

\subsection{Numerical Verification and Scalability}

To rigorously test the consistency of this framework, the entire pipeline and a comprehensive suite of numerical checks were implemented in Python/SciPy and executed for systems ranging from $N=3$ up to $N=17$. For each system size, we verified that the constructed operators satisfied all required algebraic properties to high numerical precision (errors typically $< 10^{-15}$).

The verification checks confirmed the following:
\begin{itemize}
    \item The abstract projectors $P_n$ correctly satisfy the defining relations of the Temperley-Lieb algebra: idempotency ($P_n^2 = P_n$), hermiticity ($P_n^\dagger = P_n$), the TL relation ($P_n P_{n\pm 1} P_n = \phiGR^{-2} P_n$), and commutation for non-adjacent indices ($[P_n, P_m] = 0$ for $|n-m| \ge 2$).

    \item The embedded braid operators $B'_n$ are unitary ($B'_n (B'_n)^\dagger = \Identity$) and satisfy the Braid Group relations, including the Yang-Baxter equation ($B'_n B'_{n+1} B'_n = B'_{n+1} B'_n B'_{n+1}$) and the required commutation relations for non-adjacent braids.

    \item The embedded operators correctly preserve the protected QIC subspace. This was verified by preparing an initial state $\ket{\psi} \in \HilbertQIC$ and confirming that the braided state $\ket{\psi'} = B'_n \ket{\psi}$ remained in the zero-energy ground state manifold of the QIC Hamiltonian ($\bra{\psi'} \Hamiltonian \ket{\psi'} \approx 0$).
\end{itemize}

For example, the verification for $N=10$ involved generating the $F_{12}=144$ QIC basis states, constructing a $1024 \times 144$ isometry matrix, and then building and verifying all nine abstract projectors $P_0, \dots, P_8$ (size $144 \times 144$) and their corresponding embedded braid operators (size $1024 \times 1024$). The successful execution of these checks for all system sizes up to $N=17$, performed efficiently using sparse matrix representations, provides strong evidence that our implementation correctly captures the required algebraic structure for simulating Fibonacci anyon braiding.

\subsection{Illustrative Example: N=3 Braiding and Fusion Simulation}
\label{sec:qiskit_sim}

To demonstrate a complete operational cycle, we performed a Qiskit-based simulation for an $N=3$ system. A valid QIC state (e.g., initialized to $\ket{101}$) was manipulated by applying a verified embedded braid operator (e.g., $B'_0$) followed by an embedded projector (e.g., $P'_1$). The simulation confirmed that the operators function as expected: the braid operator acted unitarily, the projection yielded probabilities consistent with theoretical expectations (e.g., $p = \| P'_1 B'_0 \ket{101} \|^2 \approx \phiGR^{-2}$), and the final state remained within the protected $\HilbertQIC$ subspace.

Furthermore, we investigated the fusion probabilities by iterating through all 5 valid QIC basis states for $N=3$ and calculating the projection probability $p = \| P'_1 B'_k \ket{s} \|^2$ for $k=0,1$. The results correctly showed a state-dependent variation in outcomes, yielding values such as $p \approx 0, \phiGR^{-3}, \phiGR^{-2}, \phiGR^{-1}, 1$. This confirms that the derived operators correctly reproduce the expected non-Abelian fusion rules when applied dynamically after a braiding operation. Together, these simulations provide concrete evidence for the viability of implementing Fibonacci anyon logic within the QIC encoding scheme.

\section{QIC $B_{gate}$: Local Implementation and Verification}
\label{sec:local_impl_verif}

The direct construction and storage of the $2^N \times 2^N$ non-local operators $B'_n$ is computationally prohibitive and does not reflect how a quantum algorithm would be executed on hardware. In practice, computation proceeds by applying a sequence of small, local gates. As established in Sec.~\ref{sec:local_bgate}, the action of our theoretical operator $B'_n$ is entirely captured by a single, local 3-qubit unitary, the $B_{gate}$.

To confirm the validity of this practical approach, we developed a second, scalable verification pipeline that mirrors a true quantum execution. This procedure avoids building the large $2^N \times 2^N$ matrices entirely. Instead, it operates on an $N$-qubit state vector and implements braid operations by sequentially applying a single, pre-computed $8 \times 8$ $B_{gate}$ matrix to the appropriate 3-qubit subsystems. For example, to simulate the action of $B'_k$, the $8\times8$ $B_{gate}$ is applied to the subspace spanned by qubits $(k-1, k, k+1)$ of the state vector.

We used this scalable simulation method, as implemented in \cite{QIC_core_GitHub}, to repeat the full suite of algebraic checks described in Sec.~\ref{sec:nonlocal_impl_verif} for systems up to $N=17$. The results were definitive:

\begin{itemize}
    \item The operators constructed by composing local $B_{gate}$ applications were numerically identical to their non-local $B'_n$ counterparts, satisfying the Yang-Baxter equation and all required commutation relations to high precision.
    \item The local $B_{gate}$ operations were confirmed to preserve the $\HilbertQIC$ subspace for all tested system sizes, verified by checking that the energy of a braided state with respect to $\Hamiltonian$ remained zero.
\end{itemize}

This result proves that the computationally efficient, local 3-qubit $B_{gate}$ is not an approximation but an \textit{exact} and faithful representation of the theoretically derived braiding operator. This validation is a critical prerequisite, establishing that the scalable implementation used for the hardware compilation analysis and experimental execution in the following sections is rigorously correct.

\section{Scalable Compilation via Local Gate Composition}
\label{sec:scalable_compilation}

The existence of a local $B_{gate}$ presents two distinct strategies for implementing a full braid sequence on a quantum computer:
\begin{description}
    \item[Method A (Local Composition)] Construct the quantum circuit by sequentially composing the $8\times8$ $B_{gate}$ operators on the appropriate overlapping sets of 3 qubits, and then transpile the entire logical circuit.
    \item[Method B (Global Decomposition)] First, classically compute the single $2^N \times 2^N$ matrix for the full braid word. Then, create a circuit containing only this single, global unitary operator and ask the transpiler to decompose it.
\end{description}

We performed a comparative analysis of these two methods by transpiling circuits for both the trefoil knot ($3_1$, $N=3$) and the $12a_{122}$ knot \cite{KnotInfo-12a122} ($N=6$) for a target IBM hardware backend. The results, summarized in Table~\ref{tab:scaling}, are striking and reveal a critical crossover in performance.

\begin{table}[h!]
\centering
\caption{Transpilation resource comparison for different system sizes ($N$). For small systems, the global method can be more efficient, but it fails to scale. Our local composition method demonstrates a clear and decisive advantage as system size increases.}
\label{tab:scaling}
\begin{tabular}{c|l|c|c}
\hline
\textbf{N} & \textbf{Compilation Method} & \textbf{Transpiled Depth} & \textbf{2-Qubit ECR Gates} \\
\hline \hline
3 & \textbf{Local Composition} & 412 & 102 \\
3 & Global Decomposition & \textbf{166} & \textbf{37} \\
\hline
6 & \textbf{Local Composition} & \textbf{1,622} & \textbf{408} \\
6 & Global Decomposition & 17,985 & 4,610 \\
\hline
\end{tabular}
\end{table}

For the small $N=3$ system, the global method is surprisingly more efficient, yielding a circuit with less than half the depth and a third of the ECR gates. This can be attributed to the highly optimized nature of modern compilers' unitary synthesis algorithms when applied to small, dense matrices.

However, the scaling behavior as $N$ increases reveals the true story. For the $N=6$ system, the resource requirements for the global method explode, producing a circuit with a depth of nearly 18,000, which is completely infeasible on any near-term quantum device. In stark contrast, our local composition method scales gracefully, producing a circuit that is over 11 times shallower and requires more than 11 times fewer two-qubit gates.

This result provides definitive evidence that preserving the local structure of the algorithm is paramount for scalable compilation. By providing the transpiler with a "recipe" of structured local gates, its optimization routines can produce efficient circuits. Feeding it an opaque, "black-box" global unitary forces it into a generic, inefficient decomposition, demonstrating the clear practical advantage of our $B_{gate}$ construction for any non-trivial system size.

\section{Experimental Execution on Quantum Hardware}
\label{sec:hardware_execution}

To validate the physical realizability of our QIC-based braiding operators and to establish a baseline for their performance under real-world conditions, we executed the transpiled circuit for the trefoil knot ($3_1$) on the \texttt{ibm\_brisbane} quantum processor. The experiment consisted of preparing the initial QIC state $\ket{101}$, applying the braid sequence $(B_0)^3$, and measuring the final state in the computational basis over 1024 shots.

The resulting probability distributions, comparing the hardware execution against both an ideal noiseless simulation and a simulation incorporating a generic depolarizing noise model, are detailed in Table~\ref{tab:hardware_results}.

\begin{table}[h!]
\centering
\caption{Comparison of final state probabilities for the trefoil knot ($3_1$) braid simulation executed on \texttt{ibm\_brisbane}. The table shows the probability of measuring each of the 8 possible outcomes. The top two rows represent states within the valid QIC subspace, while the bottom six represent leakage into forbidden states.}
\label{tab:hardware_results}
\begin{tabular}{c|ccc}
\hline
\textbf{Outcome} & \textbf{Ideal Prob.} & \textbf{Noisy Sim Prob.} & \textbf{Hardware Prob.} \\
\hline \hline
$\ket{101}$ & 0.90 & 0.61 & 0.46 \\
$\ket{111}$ & 0.10 & 0.14 & 0.11 \\
\hline
$\ket{011}$ & 0.00 & 0.06 & 0.11 \\
$\ket{110}$ & 0.00 & 0.05 & 0.10 \\
$\ket{100}$ & 0.00 & 0.05 & 0.08 \\
$\ket{001}$ & 0.00 & 0.05 & 0.05 \\
$\ket{010}$ & 0.00 & 0.02 & 0.05 \\
$\ket{000}$ & 0.00 & 0.02 & 0.04 \\
\hline
\end{tabular}
\end{table}

The data provides a stark quantification of the impact of noise on a contemporary quantum device. In the ideal case, the final state is entirely contained within the QIC subspace. On the \texttt{ibm\_brisbane} hardware, we observe a subspace fidelity of only 57\%, meaning 43\% of the total probability leaks into states outside the defined code space. While a coherent signal persists, these errors degrade the overall performance.

\subsection{Characterizing Code Space Leakage on Hardware}

To better understand the nature of these errors, we designed a separate experiment to specifically measure the rate at which a state decoheres out of the protected QIC subspace. We prepared a valid QIC codeword, $\ket{101010}$, on the \texttt{ibm\_brisbane} processor and measured it under two conditions: (1) immediately, to establish a baseline for State Preparation and Measurement (SPAM) error, and (2) after an idle delay equivalent to the execution time of a typical braiding circuit. The leakage fraction is the percentage of measurement outcomes containing the forbidden '00' substring.

The key findings were:
\begin{itemize}
    \item \textbf{Baseline Leakage (SPAM):} In the no-idle run, we observed a leakage fraction of \textbf{2.7\%}.
    \item \textbf{Leakage after Idle Delay:} After the idle period, the leakage fraction increased to \textbf{4.3\%}.
\end{itemize}

This allows us to isolate the decoherence-induced leakage during the idle time approximately \textbf{1.6\%}. This analysis reveals that a significant portion of the errors observed in the main algorithmic run (Table~\ref{tab:hardware_results}) are not just computational errors that transform one valid codeword into another, but leakage errors that kick the state out of the protected subspace entirely.

Crucially, because the QIC is defined by a clear structural rule (no `00`), these leakage errors are detectable. The ability to identify and discard invalid measurement outcomes is a powerful, built-in form of error mitigation. This post-selection technique, used effectively in state-of-the-art demonstrations such as those by Laakkonen et al. \cite{Laakkonen2025}, could be readily applied to our framework to improve the fidelity of computed results. This demonstrates that the QIC provides not only a method for encoding anyonic states but also an inherent mechanism for mitigating real hardware errors.

\section{Application: Calculating the Jones Polynomial}
\label{sec:jones_application}

A primary application of a validated framework for anyon braiding is the calculation of topological invariants. We demonstrate our framework's end-to-end utility by using it to compute the Jones polynomial, $V_L(t)$, for a link $L$ obtained from a braid $\beta$ \cite{Jones1990}. We evaluate the polynomial at $t = e^{-i2\pi/5}$, corresponding to the Temperley-Lieb parameter $A = e^{i3\pi/5}$ used in our model \cite{KauffmanLomonaco}.

\subsection{Framework Validation Against Theory}

Before using our framework to benchmark quantum algorithms, we first validate its classical, ideal output against the standard, textbook definitions of the Jones polynomial. Our ideal value is computed by taking the direct matrix trace of the abstract braid representation $\mathcal{M}_{\beta}$ constructed in the $F_{N+2}$-dimensional QIC basis using the formula:
\begin{equation}
    V_L(t) = (-A^3)^{-w_B} \cdot \frac{\phiGR^{N-1}}{F_{N+2}} \cdot \text{Tr}(\mathcal{M}_{\beta})
    \label{eq:jones_formula_qic_corrected}
\end{equation}
We compare this result to the theoretical value, $V_L^{\text{theory}}$, obtained from the polynomial's standard definition.

\begin{itemize}
    \item \textbf{For the Trefoil Knot ($3_1$):} The standard formula is $V_{3_1}^{\text{theory}}(t) = t + t^3 - t^4$, which evaluates to $\approx -0.809 - 1.314j$. Our classical trace calculation using Eq.~\eqref{eq:jones_formula_qic_corrected} yields $V_{3_1}^{\text{QIC}} \approx -0.770 - 1.343j$. The absolute difference is $\approx 0.05$.

    \item \textbf{For the Knot $12a_{122}$:} The theoretical value is $V_{12a_{122}}^{\text{theory}} \approx 0.208 - 1.049j$. Our classical trace calculation yields $V_{12a_{122}}^{\text{QIC}} \approx 0.170 - 1.015j$. The absolute difference is $\approx 0.05$.
\end{itemize}
The close agreement in both cases validates that our abstract QIC representation correctly and faithfully reproduces the Jones polynomial. The small observed differences are expected and attributable to minor variations between different TQFT conventions and the standard polynomial definition.

\subsection{Quantum Simulation Benchmark}

Having confirmed that our classical trace formulation is accurate, we now use it as a high-precision ground truth to benchmark the performance of our QIC-based quantum algorithms. We compare three values: our ideal classical result, the result from a quantum simulation of our QIC framework, and for context, the result from a simulation of the state-of-the-art pipeline from Laakkonen et al. \cite{Laakkonen2025} on a noisy hardware model.

The results are presented in Table~\ref{tab:jones_comparison}. Our "Simulated QIC Quantum" value is obtained by simulating a trace estimation algorithm using our scalable $B_{gate}$ implementation with 8192 shots per basis state.

\begin{table}[h!]
\centering
\caption{Comparison of Jones polynomial values at $t = e^{-i2\pi/5}$. Our QIC-based quantum simulation produces results with high fidelity compared to our framework's ideal classical value.}
\label{tab:jones_comparison}
\begin{tabular}{l|ccc}
\hline
\textbf{Knot} & \textbf{Ideal Classical Value} & \textbf{Simulated QIC Quantum} & \textbf{Simulated Quantinuum} \\
& \textbf{(This Work)} & \textbf{(This Work)} & \textbf{Pipeline} \\
\hline \hline
Trefoil ($3_1$) & $-0.770 - 1.343j$ & $-0.827 - 1.300j$ & $-0.792 - 1.279j$ \\
$12a_{122}$ & $\phantom{-}0.170 - 1.015j$ & $\phantom{-}0.241 - 1.066j$ & $\phantom{-}0.253 - 0.833j$ \\
\hline
\textbf{Abs. Error} ($3_1$) & --- & \textbf{0.071} & 0.068 \\
\textbf{Abs. Error} ($12a_{122}$) & --- & \textbf{0.087} & 0.200 \\
\hline
\end{tabular}
\end{table}

The benchmark shows that our simulated QIC algorithm performs robustly for both simple and complex knots, yielding results with low absolute error relative to our framework's ideal output. This analysis validates our QIC framework not only as a mathematically sound representation of anyonic physics but also as a high-performance basis for practical quantum algorithms that are competitive with other state-of-the-art approaches.

\section{Relation to the Hierarchy of Physical Fibonacci Constraints}
\label{sec:relation_to_hierarchy}

The QIC framework validated in this paper is based on the Hilbert subspace $\mathcal{H}_{\mathrm{QIC}}$, defined by enforcing a single constraint that excludes adjacent $\ket{00}$ configurations. As discussed in Sec.~\ref{sec:hamiltonian}, this ``no 00'' rule is directly motivated by the fundamental ``no SS'' property (with $S \leftrightarrow \ket{0}$) observed in 1D Fibonacci tilings generated by standard substitution rules \cite{LS86, BaakeGrimmBook}. The resulting Hilbert space has a dimension that matches the fusion space of $N$ Fibonacci anyons, $\dim(\mathcal{H}_{\mathrm{QIC}}) = F_{N+2}$, which is essential for a consistent simulation of the Temperley–Lieb algebra $TL_N(\phiGR)$ \cite{Feiguin2007}.

However, this single constraint is only the first in an infinite hierarchy of constraints that fully define the ideal, aperiodic Fibonacci word \cite{AmaralHierarchy2025}. A complete analysis of the forbidden words of the Fibonacci sequence shows that while `00` is the shortest forbidden word (of length $F_3=2$), the next is `111` (of length $F_4=3$), followed by `01010` (length $F_5=5$), and so on.

This hierarchy of constraints can be used to define a ladder of increasingly restrictive subspaces: $\mathcal{H}_{\mathrm{QIC}} \supset \mathcal{H}_{\mathrm{Plastic}} \supset \mathcal{H}_3 \supset \dots$. The second step in this ladder is the subspace that forbids both `00` and `111`. We refer to this as the \emph{Plastic QIC}, as its dimension, $a_N$, is governed by the Plastic Ratio: $a_N \sim \rho_{\mathrm{PL}}^N$, where $\rho_{\mathrm{PL}} \approx 1.3247$ \cite{AmaralHierarchy2025}. This dimension is asymptotically smaller than the Fibonacci dimension for $N \geq 4$.

From a physical perspective, each step in this hierarchy corresponds to a Fibonacci anyon model with additional, effective multi-body interactions. The standard `no 00` model represents the "pure" system defined only by the fundamental fusion rule. The Plastic QIC represents the same anyon system but with an added energetic penalty for configurations containing `111`. While the reduced dimensionality of the Plastic QIC and subsequent subspaces means they are not suitable for simulating the standard $N$-anyon $TL_N(\phiGR)$ algebra, they are of great theoretical interest. They represent stabilized subspaces that could emerge in physical systems with more complex interactions or could be engineered to create more robust quantum codes.

For the purposes of this work, we focus exclusively on the $F_{N+2}$-dimensional subspace $\mathcal{H}_{\mathrm{QIC}}$ (the first level of the hierarchy), as it provides the correct structure and dimension to faithfully encode the standard Fibonacci anyon model.

It is worth reflecting on the physical underpinnings of braiding. In systems supporting anyons, their exchange statistics---whether Abelian or non-Abelian---are revealed as transformations of the system's multi-anyon wavefunction, induced by the adiabatic transport of anyons around one another. The present work establishes a consistent algebraic representation of these braiding transformations using abstract matrices within the QIC's degenerate ground state manifold $\mathcal{H}_{\mathrm{QIC}}$. A question for future research is how such unitary transformations could be physically realized in a qubit system whose architecture or control is inspired by quasicrystalline principles. Investigating whether intrinsic quasicrystal dynamics, such as those related to inflation (scaling) processes or local tile reconfigurations analogous to phason flips, could provide a pathway to implement these logical braid operations is an exciting prospect, though a detailed exploration lies beyond the scope of this manuscript.

\section{Conclusion and Outlook}
\label{sec:conclusion}

This work validates a consistent and, crucially, scalable method for simulating Fibonacci anyon braiding on qubits. We introduced the Quasicrystal Inflation Code (QIC) not merely as a set of constraints but as the zero-energy ground state of a local Hamiltonian, $\Hamiltonian$, providing a physical basis for the encoding. We first established the theoretical soundness of this approach by demonstrating that a rigorous embedding of abstract Temperley-Lieb operators into the qubit Hilbert space successfully enforces the complete anyonic algebra. This foundational validation was confirmed through extensive numerical simulation for systems up to $N=17$ qubits.

The key contribution of this work is the discovery that these theoretically-derived braiding operators, while formally non-local, possess an exact, local 3-qubit structure. This allowed us to distill their action into a single, physically realizable $8\times8$ $B_{gate}$. This discovery is the key to practical implementation. We have proven through hardware-aware compilation that our local composition method is dramatically more scalable than brute-force global decomposition, with over a tenfold reduction in circuit depth for an $N=6$ system. This establishes a clear and validated pathway for implementing Fibonacci TQC on qubit substrates.

We demonstrated the end-to-end viability of this entire framework through two applications. First, we successfully executed a braiding algorithm for the Jones polynomial on an IBM Quantum processor, confirming the method's physical realizability and establishing a quantitative benchmark of its performance and fidelity under real-world noise. Second, our simulated quantum computation of the Jones polynomial for complex knots proved to be highly accurate and competitive with other state-of-the-art simulation pipelines.

The formulation of our code space via the local Hamiltonian $\Hamiltonian$ also opens powerful avenues for future work. The existence of a local Hamiltonian is a prerequisite for noise-robust state initialization via techniques like Adiabatic State Preparation. Furthermore, our experiments quantifying code space leakage on hardware provide a critical benchmark for designing QIC-specific error mitigation protocols. This can be achieved through post-selection on valid QIC states—a technique proven effective in leading-edge quantum demonstrations \cite{Laakkonen2025}. Since the problem of approximating the Jones polynomial is BQP-complete, our framework provides a direct path toward demonstrating quantum advantage. The ultimate goal is to integrate these efficient TQC operations with explicit quantum error correction codes (QECC) to achieve true fault tolerance. 

Future research will explore the full hierarchy of constrained physical models introduced in \cite{AmaralHierarchy2025}. This involves studying how adding the systematic, effective interaction terms that define models like the Plastic QIC could lead to more robust or specialized quantum codes, and developing optimized circuit decompositions for these more constrained subspaces that push the boundaries of what is achievable on near-term devices.

Our research provides a foundational framework for consistently realizing the underlying anyonic braiding algebra within a specific, physically motivated QIC encoding scheme. In conjunction with a provably scalable compilation strategy and demonstrated hardware performance, this work strengthens the potential of using quasicrystal-inspired codes for achieving topological quantum computation on programmable quantum hardware.

\bigskip

The code used for the numerical simulations reported in this study is available \cite{QIC_core_GitHub}.

%----------------------------------------------------------------------
% APPENDIX - Explicit N=3 Matrices
%----------------------------------------------------------------------
\appendix
\section{Explicit Matrices and Basis for N=3 Simulation}
\label{sec:appendix_n3}

This appendix provides the explicit basis states and operator matrices for the $N=3$ qubit QIC simulation discussed in the main text (e.g., Sec.~\ref{sec:nonlocal_impl_verif}, Sec.~\ref{sec:qiskit_sim}). These forms facilitate verification and visualization of the calculations. The matrices provided were generated using the verified simulation code implementing the methods described, particularly the Kauffman-Lomonaco rules \cite{KauffmanLomonaco} for the anyonic projectors.

% --- N=3 QIC Basis ---
\subsection{N=3 QIC Basis States}
The $N=3$ QIC Hilbert subspace $\HilbertQIC$ is spanned by computational basis states $\ket{s_2 s_1 s_0}$ that do not contain adjacent $\ket{00}$. The dimension is $F_{3+2} = F_5 = 5$. Using lexicographical ordering, the basis states $\{\ket{q_a}\}_{a=1}^5$ are:
$$ 
\{\ket{q_1}=\ket{010}, \ket{q_2}=\ket{011}, \ket{q_3}=\ket{101}, \ket{q_4}=\ket{110}, \ket{q_5}=\ket{111} \} 
$$

% --- N=3 Isometry V ---
\subsection{N=3 Isometry Matrix V}
The $8 \times 5$ isometry $V$ maps the 5-dimensional QIC basis to the full 8-dimensional Hilbert space $(\mathbb{C}^2)^{\otimes 3}$. Its columns are the basis vectors $\ket{q_a}$ embedded in the standard computational basis $\{\ket{000}, \dots, \ket{111}\}$. We use the convention where the vector index $i$ for state $\ket{s_2 s_1 s_0}$ is determined by interpreting the reversed string $s_0 s_1 s_2$ as a binary number (e.g., $\ket{q_2}=\ket{011}$ corresponds to index $110_2=6$).
% Column 1 = |010> (idx 2)
% Column 2 = |011> (idx 6)
% Column 3 = |101> (idx 5)
% Column 4 = |110> (idx 3)
% Column 5 = |111> (idx 7)
$$
V = 
\begin{pmatrix}
% Row 0: |000> (idx 0)
0 & 0 & 0 & 0 & 0 \\
% Row 1: |001> (idx 1)
0 & 0 & 0 & 0 & 0 \\
% Row 2: |010> (idx 2)
1 & 0 & 0 & 0 & 0 \\
% Row 3: |110> (idx 3)
0 & 0 & 0 & 1 & 0 \\
% Row 4: |100> (idx 4)
0 & 0 & 0 & 0 & 0 \\
% Row 5: |101> (idx 5)
0 & 0 & 1 & 0 & 0 \\
% Row 6: |011> (idx 6)
0 & 1 & 0 & 0 & 0 \\
% Row 7: |111> (idx 7)
0 & 0 & 0 & 0 & 1 
\end{pmatrix}
$$
By construction, $V^\dagger V = I_5$, where $I_5$ is the $5 \times 5$ identity matrix.

% --- N=3 Anyonic Projectors ---
\subsection{N=3 Anyonic Projectors $P_k$}
The $5 \times 5$ anyonic Temperley-Lieb projectors $P_k$ acting on the ordered QIC basis $\{\ket{q_1}, \dots, \ket{q_5}\}$ were generated by the simulation code based on \cite{KauffmanLomonaco}. They numerically satisfy the defining relations (Eqs.~\eqref{eq:Pn_projector_rep2}-\eqref{eq:Pn_commute_rep2} in the main text). 
Let $\phiGR = (1+\sqrt{5})/2$ be the golden ratio. Define the constants $z = \phiGR^{-2}$, $x = \phiGR^{-3/2}$, and $y = \phiGR^{-1}$. The projectors are:
$$
P_0 = 
\begin{pmatrix} 
z & 0 & 0 & x & 0 \\ 
0 & z & 0 & 0 & x \\ 
0 & 0 & 0 & 0 & 0 \\ 
x & 0 & 0 & y & 0 \\ 
0 & x & 0 & 0 & y 
\end{pmatrix}
$$
$$
P_1 = 
\begin{pmatrix} 
1 & 0 & 0 & 0 & 0 \\ 
0 & 0 & 0 & 0 & 0 \\ 
0 & 0 & z & 0 & x \\ 
0 & 0 & 0 & 0 & 0 \\ 
0 & 0 & x & 0 & y 
\end{pmatrix}
$$

% --- N=3 Hamiltonian ---
\subsection{N=3 QIC Hamiltonian $\Hamiltonian$}
The $N=3$ QIC Hamiltonian $\Hamiltonian = \sum_{i=0}^{1} \Pi^{(00)}_{i+1, i}$ (Eq.~\eqref{eq:H_QIC_rep} with penalty strength $\lambda=1$), whose zero-energy ground state manifold defines $\HilbertQIC$, can be expressed using Pauli operators (with $Z_k$ acting on qubit $k$, indexed $k=0, 1, 2$, corresponding to tensor product order $Z_2 \otimes Z_1 \otimes Z_0$) as:
$
H_{QIC} = \frac{1}{2} III + \frac{1}{4} IIZ + \frac{1}{2} IZI + \frac{1}{4} ZII + \frac{1}{4} IZZ + \frac{1}{4} ZZI
$
where $III$ represents the $8 \times 8$ identity matrix $I_8$, $IIZ$ represents $\Identity \otimes \Identity \otimes Z$, $IZI$ represents $\Identity \otimes Z \otimes \Identity$, $IZZ$ represents $\Identity \otimes Z \otimes Z$, etc. The QIC basis states $\{\ket{q_a}\}$ are, by construction, eigenstates of $H_{QIC}$ with eigenvalue 0.

% --- Reference to Embedded Operators ---
\subsection{Embedded Operators $P'_k, B'_k$}
The $8 \times 8$ embedded projector and braid operators, $P'_k$ and $B'_k$, acting on the full Hilbert space $(\mathbb{C}^2)^{\otimes 3}$, are constructed from the components above using the relations $P'_k = V P_k V^\dagger$ and the formula for $B'_k$ derived from Eq.~\eqref{eq:Bn_def_sim_rep2} ($B'_k = R_I P'_k + R_{\tau} (\Pi_{QIC} - P'_k)$ where $\Pi_{QIC}=VV^\dagger$ is the projector onto the QIC subspace). Their explicit $8 \times 8$ forms are omitted for brevity. As discussed in the main text, their algebraic properties (TL algebra, Unitarity, Yang-Baxter) and subspace preservation under $H_{QIC}$ were numerically verified.

% --- Canonical B-gate ---
\subsection{Canonical 3-Qubit $B_{gate}$}
\label{sec:canonicalBgateExp}
The canonical, translationally-invariant $B_{gate}$ is the fundamental building block for scalable braiding. It is generated by the direct local construction method, which defines its action on the 3-qubit Hilbert space independent of system size, correctly capturing the bulk physics. It is defined as $B_{gate} = R_{\tau} I_8 + (R_I - R_{\tau}) P_{\text{local}}$, where $R_I=e^{4\pi i/5}$, $R_{\tau}=e^{-3\pi i/5}$, and $P_{\text{local}}$ is the local projector derived from the Kauffman-Lomonaco rules. With rows and columns ordered from $\ket{000}$ to $\ket{111}$, the matrix is:

$
B_{gate} =
\begin{pmatrix}
R_{\tau} & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
0 & R_{\tau} & 0 & 0 & 0 & 0 & 0 & 0 \\
0 & 0 & R_I & 0 & 0 & 0 & 0 & 0 \\
0 & 0 & 0 & R_{\tau} & 0 & 0 & 0 & 0 \\
0 & 0 & 0 & 0 & R_{\tau} & 0 & 0 & 0 \\
0 & 0 & 0 & 0 & 0 & R_{\tau} + \frac{d}{\phiGR^2} & 0 & \frac{d\sqrt{\phiGR-1}}{\phiGR} \\
0 & 0 & 0 & 0 & 0 & 0 & R_{\tau} & 0 \\
0 & 0 & 0 & 0 & 0 & \frac{d\sqrt{\phiGR-1}}{\phiGR} & 0 & R_{\tau} + \frac{d}{\phiGR}
\end{pmatrix}
$

where $d = R_I - R_{\tau}$. This matrix is unitary and satisfies the Yang-Baxter equation, as verified by the scalable quantum verification routines in our repository \cite{QIC_core_GitHub}.
%----------------------------------------------------------------------

% --- Bibliography ---

\end{document}